Reports 1-1 of 1 Clear search Modify search
Injection system (Pre-commissioning)
chiummo, Spinicelli - 21:52 Thursday 17 September 2020 (49442) Print this report
INJ recovery after DAQ work

After DAQ team managed to give us back the signals for the Injection system, we started to recover the normal state. We enabled the control of the EIB, locked the PMC, managed to close the BPC loop and reduced the related corrections by acting on the usual M6-M8 picomotors, then aligned and locked the IMC cavity.

At some point, while trying to progress with automatic alignment (more on this soon), I (AC) deemed wise to discharge the integrator of the MC DSP card, as done several times in the past. Unfortunately, this time it resulted in triggering the guardian, but also - for reasons to me unknown - the impossibility to communicate anymore with MC suspension. The work is therefore suspended until tomorrow morning.

To the benefit of the DAQ team check, it is also to be noticed that we managed also to lock for some minutes the reference cavity.

We leave the PMC in scanning mode, BPC with static corrections, and the beam blocked on the Laser bench.



Comments to this report:
casanueva, chiummo, melo, spinicelli, vardaro - 19:30 Saturday 19 September 2020 (49449) Print this report

INJ recovery continued on Friday, after Thursday problems with DSP card. The final status is (details later here): INJ locked from Friday evening, RFC locked, PSTAB working, Fmoderr tuned and AA in drift control. The system appears to be stable overnight with the exception of three short fast unlock (see fig.1). On Monday, we will test the automatic alignment in full bandwidth.

More details:
After the problem with MC suspensions we started back the alignment by closing BPC loop. We then looked for a MC_F0_Z position close to RFC TEM00 and tuned fmoderr to 0. Checking the RFC phase, we remarked a strong phase shift of the error signal of the RFC (fig.4) then we correctly tuned it (fig. 5).
Then, recovering the position of IB bench and MC to last alignment in which AA was working (49402) brought the system in an unstable state  (i.e. Tx/Ty of both MC and IB showed wide oscillation). Those oscillations were probably preventing AA system to work properly. We decided thus to move toward another previous position of IMC alignment by tuning directly the beam position on the MC_Cam (see fig. 2 and 3). Once the position finely tuned maximizing IMC_TRA and optimizing reflection, oscillations in Tx/Ty (in particular Tx for MC) were strongly reduced. Finally, we manually brought to 0 in  ACL the offsets in tx/ty/tz error signal for both IB and MC.
After that, we measured again the sensing matrix with AA and RFC loops open, by injecting noise (a line at 2 Hz) on each AA DoF (see table below):

MC ty 1284484080.00
IB ty 1284483757.00
MC tx 1284483974.00
IB tx 1284483680.00
IB tz 1284483835.00


In addition, to test the correctness of the new matrix, we individually tested again each degree of freedom by injecting noise, and verified that only the relative reconstructed error signal was acting.

We left AA loop in drift control.


- check offsets on galvos
- check phase modulation of QD FF/NF
- close IBz loop
- close AA loop on full bandwidth





Images attached to this comment
chiummo, melo, spinicelli - 18:22 Monday 21 September 2020 (49461) Print this report

After the today problem with the network and the slave diodes we recovered the alignment obtained on Friday. To be noted that before the signal interruption, IMC stayed locked in drift control and was stable during the whole weekend without major issues (see fig. 1).

We then closed IBz loop and finally close the AA in full bandwidth. The system appears to be stable, we will check the behavior overnight.

Images attached to this comment
Search Help